# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

### v1

@@ Abstract
The shuffle model of differential privacy (DP) has recently emerged as a powerful one for decentralized computation without fully trustable parties. Since it anonymizes and permutes messages from clients through a shuffler, the privacy can be amplified and utility can be improved. However, the shuffling procedure in turn restricts its applications only to statistical tasks that are permutation-invariant.
This work explores the feasibility of shuffle privacy amplification for prevalent non-statistical computations: spatial crowdsourcing, combinatorial optimization, location-based social systems, and federated learning with incentives, which suffer either computationally intractability or intolerable utility loss in existing approaches (e.g., secure MPC and local DP). We proposes a new paradigm of shuffle model that can provide critical security functionalities like message authorization and result access control pertained to permutation-equivariant tasks, meanwhile maintaining the most of privacy amplification effects. It incurs almost the same computation/communication costs on the client side as the non-private setting, and permits the server to run arbitrary algorithms on (noisy) client information in plaintext. Our novel technique is introducing statistically random identity into DP and force identical random distribution on all clients, so as to support secure functionalities even after message shuffling and to maintain privacy amplification simultaneously. Given that existing DP randomizers fails in the new shuffle model, we also propose a new mechanism and prove its optimality therein. Experimental results on spatial crowdsourcing, location-based social system, and federated learning with incentives, show that our paradigm and mechanism is fast as non-private settings, while reducing up to $90\%$ error and increasing utility performance indicates by $100\%$-$300\%$ relatively, and can be practical under reasonable privacy budget.
@@
differential privacy, shuffle model, digital identities, spatial crowdsourcing, federated learning
@@ I Introduction
Personal information fuels a wide array of data-driven applications, e.g. statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, mobile crowdsensing, combinatorial optimization, and location-based social systems. These applications deliver substantial value but depend on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse. Data privacy concerns are escalating, especially after several high-profile data breach incidents. Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation, the California Consumer Privacy Act, and China’s Personal Information Protection Law, many users still distrust service providers and are hesitant to consent to the use of their data. To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing without relying on a trusted third party.
@@
Two prominent technologies addressing this problem are secure multiparty computation (MPC) [88] and differential privacy (DP) [28]. MPC uses interactive cryptographic protocols to allow mutually untrusted parties to jointly evaluate a function on their private data, with each party receiving an output but nothing else. However, the cryptographic nature of MPC imposes heavy overheads, making it challenging to scale for practical use. DP, on the other hand, adds random noise to data to ensure privacy, making it computationally more feasible since the computation is carried out on sanitized data without heavy cryptographic machinery. Traditional central DP still requires a trusted curator [28], but the local model (LDP [53]) allows each user to sanitize their data locally before sending it out. Due to its practicality and low trust assumption, LDP has been adopted by companies like Apple [74], Microsoft [24], and Google [30] in their real-world systems. However, LDP’s drawback is that each user must independently inject sufficient random noise to guarantee privacy, leading to significant utility loss.
@@
Recently, within the realm of DP research, the shuffle model [17, 29] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to the computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [31, 32]. Consequently, to achieve a predefined global privacy goal, each user consumes less privacy budget when sanitizing their data locally, meaning less noise needs to be added. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [10, 37, 38, 38, 39].
@@
At first glance, private individual computation seems unattainable within the shuffle model because the need for individualized output clashes with the anonymization required for privacy amplification. However, this is not necessarily true. Our observation is that many individual computing functionalities are equivariant to shuffling. That is, the permutation applied to the inputs does not affect the computation. This means shuffling does not prevent the server from producing personalized answers for each client—the server does not know and does not need to know which answer is for whom. Nevertheless, there is a problem: how to pass the output back to the right user without breaking anonymity. A straightforward approach is to let the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [41]), but this method is costly (due to storing communication states) and may be subject to de-anonymization attacks on anonymous channels [63, 65]. Additionally, the duplex connection is not supported in current shuffler implementations, such as the one in the seminal work [17]. This is the first technical challenge we need to address.
@@
The second challenge we face is how to design optimal randomizers for individual computation in the shuffle model. Often the randomizer in a shuffle model protocol needs to be tailored to the tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user (instead of the statistical accuracy of the population). This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for the new setting. Therefore we need to take a fresh look at the fundamental privacy-utility trade-offs and re-design the underlying randomizers.
In this work, we introduce a new paradigm of the shuffle model that provides security functionalities (e.g., message authorization, integrity verification, and access control) for permutation-equivariant computations, and term it the private individual computation model. The properties of the model are as follows.
  • I.
    Security: The server and other clients can verify the owner of a message using a fingerprint submitted alongside the message. The server can also securely deliver personalized backward information to each owner (i.e., access control). These functionalities enable the secure completion of equivariant tasks.
  • II.
    Privacy: There is no degradation of local privacy while providing these additional security functionalities. Furthermore, when task results are sparse (e.g., each user is matched to one taxi driver in spatial crowdsourcing), the degradation of shuffle privacy amplification is negligible. Additionally, our local and shuffle privacy guarantees are information-theoretic (as long as the shuffling is statistically random). This distinguishes our new model from computational DP [62], which also combines cryptographic functions and privacy but guarantees differential privacy only against polynomial-power adversaries.
  • III.
    Utility: Since clients still benefit from privacy amplification via shuffling, the utility is significantly improved. With the increasing of number of clients, the utility increases and approaches to non-private settings. In addition, the server in our model can run arbitrary algorithms on the (noisy) messages in plaintext. Our model is compatible with existing achievements and future advancements in server-side algorithms (e.g., noisy-aware matching algorithms for spatial crowdsourcing).
  • IV.
    Scalability: As the server runs algorithms on plain-text messages and there is no heavy computation burden on the client side, our model is highly efficient and scales to systems with massive clients. This eliminates the need for secure multi-party computation which necessitates impractical costs. In our model, the computation/communication costs are essentially the same to non-private settings (see Appendix -I).
Main techniques. Our approach to realizing additional security functionalities is based on two critical observations: (1) If the message payload satisfies $\epsilon$-LDP, appending arbitrary random variables (independent to the private value) to the payload incurs no additional privacy loss. Specifically, when random variables are sampled from an identical distribution for all clients, there is no excess loss in privacy amplification via shuffling. (2) Post-processing the message payload and appended random variables does not incur excess privacy loss. The first observation motivates us to append a random fingerprint (e.g., a public key) to the payload, allowing the client to use the corresponding secret key to negotiate shared secret and establish secure communication with other parties in later phases. A side consequence is that the server can use the public key pertained to each message to securely deliver information to the message owner. The second observation motivates us to further append a signature to protect the integrity of the fingerprint and payload.
Challenge (2): Optimal randomizers for non-statistical computations in the shuffle model. In the local model of DP, existing research [83, 66, 77, 81, 82] primarily focuses on designing randomizers in a high privacy regime (e.g., local budget $\epsilon=O(1)$). Conversely, in the shuffle model, the local budget can extend to $\tilde{O}(\log(n))$. For statistical tasks within this model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of non-statistical computation tasks in the shuffle model—often involving multi-dimensional user data and emphasizing an individual report’s accuracy (instead of statistical accuracy of the population)—has not been explored. This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for non-statistical computation in the shuffle model. The fundamental privacy-utility trade-offs in non-statistical computation for this new model also present significant theoretical interest.
Our work establishes the error lower bounds for single reports in non-statistical computation within the shuffle model, demonstrating the sub-optimality of existing randomizers. To address this, we introduce a novel series of local randomizers—the Minkowski response—aligned with these error lower bounds. The core principle involves reporting values within a calibrated Minkowski distance from the true value with high probability.
The main contributions of this work are summarized as follows:
  • We propose equivariant shuffle model of DP, which supports various security functionalities and benefit large-scale equivariant computations from privacy amplification. The new model exhibits almost the same computation/communication costs on client/server side as non-private settings.
  • We establish tight minimax error lower bounds of $\tilde{O}((n\epsilon^{2}_{c})^{-\frac{2}{d+2}})$ for the new model of DP, where $d$ is the user data dimension, $\epsilon_{c}$ is global privacy budget, and $n$ is the number of users. Since existing data randomizers fail, we then design an attainable randomizer: Minkowski response, for the new model to match the lower bound.
  • Through experiments on three typical permutation-equivariant tasks: privacy-preserving spatial crowdsourcing, location-based nearest neighbor queries, and federated learning with incentives, we demonstrate that the new model, in conjunction with proposed randomizers, reduces error by $90\%$ meanwhile incurring negligible excess burdens in both the client and server side.
@@
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC model. Section VI provides optimal LDP randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
@@ II Related Works
This section reviews various approaches to private computation, primarily concentrating on permutation-equivariant tasks.
II-A Cryptography Methods
Numerous works aim to securely solve matching and allocation problems involving personal sensitive information from multiple parties, revealing only the output. They utilize homomorphic encryption [2, 44], secret sharing [3] or garbled circuits (e.g., in [40, 15]) to perform secure multi-party computations (MPC). Specifically, when there is a trusted message shuffler, every client can establish information-theoretically secure communication with other clients and accomplish arbitrary multi-party computation (including DP algorithms) within a few rounds [48, 13]. In general, MPC-based approaches impose high computation and interaction overheads, especially when the computation algorithm gets sophisticated. As comparison, the proposed paradigm in this work permits running arbitrary algorithms on plaintext, and requires only semi-trustness/honest-but-curious assumption on the shuffler (as the shuffler only sees ciphertexts, see Section V).
@@ II-C Shuffle Model of DP
The recently proposed shuffle model [17, 29] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust). Depending on the number of messages each client can send to the intermediate shuffler, the shuffle model can be categorized as single-message [9, 29, 32] and multi-message [10, 36]. The single-message shuffle model leverages privacy amplification via shuffling to enhance data utility compared to the local model. A substantial body of work [29, 31, 32] demonstrates that $n$ shuffled messages from clients, each adopting a same $\epsilon$-LDP randomizer, can actually preserve $\tilde{O}(\sqrt{e^{\epsilon}/n})$-DP. By removing the constraint of sending one message, the multi-message shuffle model can achieve better utility than the single-message model and might be comparable to the curator model (e.g., in [10, 36]). However, each multi-message protocol is tailored to a specific statistical query (e.g., summation), rendering them unsuitable for permutation-equivariant tasks with non-linear computations. There is a line of works on the shuffle model for private information retrieval (e.g., in [48, 47, 34] with cryptography security and in [78, 1] with statistical DP), where the query is represented as multiple secret shares before sent to the shuffler, and the server holding the database entries returns linear-transformed entries for each query, using the duplex shuffled communication channel. This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [63, 65], and is pertained to the linear computation in private information retrieval. It can not be extended to other PIC tasks that involve with non-linear computations, and can not handle party-to-party communication in tasks like spatial crowdsourcing and social systems.
Overall, existing works in the shuffle model primarily focus on statistical queries that are naturally permutation-invariant. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms). Technically, our work extends the shuffle model to support identity-related security primitives, such as message authentication and access control.
@@ II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multi-party computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [43]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [62] against polynomial-time adversaries. Computational DP protocols inherited the computation/communication complexity of MPC (refer to garbled-circuits-based approaches in Table VIII).
In contrast to enhancing relaxed privacy in the cryptographic cipher domain, our work initializes the study in the opposite direction, by integrating simple security functionalities into the privacy realm. Our model exposes intermediate results with information-theoretical DP guarantee. Therefore, it is significantly more efficient (i.e., without computation on ciphertext and requiring few rounds of interaction) and flexible (i.e., supporting arbitrary computations on noisy plaintext) than MPC with computational DP.
III Background and Problem Settings

### v2

# Beyond Statistical Estimation: Differentially Private Individual Computation via Shuffling

@@ Abstract
In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge. Technologies like Differential Privacy (DP) have been pivotal in addressing these concerns. The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling. These benefits have led to significant interest in the shuffle model. However, the computation tasks in the shuffle model are limited to statistical estimation, making the shuffle model inapplicable to real-world scenarios in which each user requires a personalized output. This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations. PIC enables personalized outputs while preserving privacy, and enjoys privacy amplification through shuffling. We propose a concrete protocol that realizes PIC. By using one-time public keys, our protocol enables users to receive their outputs without compromising anonymity, which is essential for privacy amplification. Additionally, we present an optimal randomizer, the Minkowski Response, designed for the PIC model to enhance utility. We formally prove the security and privacy properties of the PIC protocol. Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
@@
Differential Privacy, Private Individual Computation, Shuffle Model, Privacy-Utility Trade-offs
@@ I Introduction
Personal information fuels a wide array of data-driven applications, e.g. statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, e-health, social networks, and smart cities. These applications deliver substantial value but rely on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse. Data privacy concerns are escalating, especially after several high-profile data breach incidents. Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation and the California Consumer Privacy Act, many users still distrust service providers and are hesitant to consent to the use of their data. To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing.
@@
Two prominent technologies addressing this problem are secure multiparty computation (MPC) [91] and differential privacy (DP) [31]. MPC employs interactive cryptographic protocols that enable mutually untrusted parties to jointly compute a function on their private data, ensuring each party receives an output while learning nothing else. Despite its strong privacy guarantees, the cryptographic nature of MPC results in substantial overhead, posing scalability challenges for practical applications. In contrast, DP enhances privacy by adding random noise to the data, allowing computation to be performed on sanitized data without the need for heavy cryptographic machinery. Traditional central DP relies on a trusted curator [31], but the local model (LDP [54]) enables users to sanitize their data locally before sharing it. LDP’s practicality and minimal trust requirements have led to its adoption by companies such as Apple [77], Microsoft [27], and Google [33] in their real-world systems. However, LDP’s primary drawback is that each user must independently add sufficient random noise to ensure privacy, which can significantly impact the utility of the data.
@@
Recently, within the realm of DP research, the shuffle model [13, 32] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to a computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [34, 35]. Consequently, to achieve a predefined global privacy goal, less noise must be added when users sanitize their data locally. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [8, 40, 41, 41, 42].
@@
At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification. However, this is not necessarily the case. Our observation is that many individual computing tasks are equivariant to shuffling, meaning that the permutation applied to the inputs does not affect the computation. Therefore, shuffling does not prevent the server from producing personalized answers for each client – the server does not need to know which answer belongs to whom. Nevertheless, there is a challenge: how to return the output to the correct user without compromising anonymity. One straightforward approach is to have the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [44]). However, this method is costly due to the need to store communication states and may be vulnerable to de-anonymization attacks on anonymous channels [64, 66]. Additionally, current shuffler implementations, e.g. the one described in the seminal work [13], do not support such duplex connections. Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which is intractable in the duplex shuffle channel. Addressing this issue is the first technical challenge we need to overcome.
@@
The second challenge we face is designing optimal randomizers for individual computation within the shuffle model. Often, the randomizer in a shuffle model protocol should be tailored to specific tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as demonstrated in histogram estimation [39, 34] and one-dimensional summation estimation [7]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population. This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [88]), as well as prevalent randomizers (such as adding Laplace [31] or Staircase [38] noises), less effective for the new setting. Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
Our contributions In this work, we introduce a new paradigm extending the shuffle model that allows a wider range of permutation-equivariant tasks to be computed with DP guarantee and can enjoy privacy amplification provided by shuffling. We term the new paradigm Private Individual Computation (PIC in short). We define PIC formally as an ideal functionality, which captures its functional and security properties. We also provide a concrete protocol, with formal proof, that can realize the ideal functionality.
Similar to the shuffle-DP protocols, each user adds noise locally to their data and encapsulates it (and possibly other auxiliary data) into an encrypted message under the public key of the computation server. The shuffler then shuffles the encrypted messages, before sending them to the computation server. The server can decrypt these messages and perform the permutation-equivariant computation. What differs is that each user also includes a one-time public key within the encrypted message. This one-time public key serves two purposes: (1) it allows the server to encrypt the computation result such that it can only be decrypted by the owner of the corresponding private key, and (2) it acts as a pseudonym for the key owner. This approach addresses our first challenge with minimal overhead: the server can publish a list where each entry consists of a public key along with the computation result encrypted under this key. Users can download the entire list and decrypt the entry associated with their own public key, maintaining their anonymity. Additionally, users can establish secure communication channels with other matched parties using their public keys to eventually complete PIC tasks.
Another main contribution is the development of an asymptotically optimal randomizer specifically designed for the PIC model. This randomizer is based on an LDP mechanism we call Minkowski Response. The primary goal of this design is to enhance utility, measured by the single-report error, which is the expected squared error between a user’s true data value and its sanitized version. To achieve this, a Minkowski distance $r$ is determined based on the privacy budget. The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$. The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere in the entire output domain with a relatively low probability.
We formally prove the security and privacy properties of the PIC protocol. Additionally, we provide a theoretical analysis of the utility bounds achievable by protocols in the PIC model and the Minkowski Response mechanism. Our analysis demonstrates that asymptotically, the error upper bound of the Minkowski Response matches the error lower bound for all possible randomizers in the PIC model, thereby achieving optimality. Alongside theoretical analysis, we conducted extensive experiments using real-world applications and datasets. The evaluation confirms that computations conducted in the PIC model exhibit significantly better utility than those in the LDP model. Furthermore, the performance of the Minkowski randomizer, measured by single-report error and task-specific utility metrics, surpasses that of existing LDP randomizers commonly used in the shuffle model.
@@
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC protocol. Section VI provides optimal user-side randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
@@ II Related Works
This section reviews various approaches to private computation, primarily concentrating on non-statistical tasks.
II-A Secure Multiparty Computation
Secure Multiparty Computation (MPC) is a fundamental cryptographic primitive that enables multiple parties to jointly compute a function over their inputs while no party learns anything beyond their own input and the final output of the computation. MPC was first conceptualized by Andrew Yao in the 1980s, and it has been proven that any computable function can be realized by MPC [90]. MPC relies on cryptographic protocols that exchange encrypted messages among parties. To allow computing on encrypted data, primitives such as homomorphic encryption [67, 19], secret sharing [14, 26], or garbled circuits [91] can be employed. The primary challenge in MPC lies in balancing privacy and efficiency. While MPC offers strong privacy guarantees, it often suffers from significant computational and communication overheads, which makes scaling to large datasets or numerous parties difficult. Recent research in MPC has been focusing on optimizing protocols and practical implementations [23, 55, 16, 24, 70, 25, 75, 15]. Despite a significant improvement, MPC still faces efficiency issues that hinder its widespread real-world deployment.
@@ II-C Shuffle Model of DP
The recently proposed shuffle model [13, 32] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust). Depending on the number of messages each client can send to the intermediate shuffler, the shuffle model can be categorized as single-message [7, 32, 35] and multi-message [8, 39]. The single-message shuffle model leverages privacy amplification via shuffling to enhance data utility compared to the local model. A substantial body of work [32, 34, 35] demonstrates that $n$ shuffled messages from clients, each adopting a same $\epsilon$-LDP randomizer, can actually preserve $\tilde{O}(\sqrt{e^{\epsilon}/n})$-DP. By removing the constraint of sending one message, the multi-message shuffle model can achieve better utility than the single-message model and might be comparable to the curator model (e.g., in [8, 39]). However, each multi-message protocol is tailored to a specific statistical query (e.g., summation), rendering them unsuitable for permutation-equivariant tasks with non-linear computations. There is a line of works on the shuffle model for private information retrieval (e.g., in [49, 48, 37] with cryptography security and in [81, 1] with statistical DP), where the query is represented as multiple secret shares before sent to the shuffler, and the server holding the database entries returns linear-transformed entries for each query, using the duplex shuffled communication channel. This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [64, 66], and is pertained to the linear computation in private information retrieval. It can not be applied to other PIC tasks (such as combinatorial optimization and federated learning with incentive) that involve with non-linear computations, and can not provide secure user-to-user communication needed in tasks like spatial crowdsourcing and social systems.
Overall, existing works in the shuffle model primarily focus on statistical queries. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms).
@@ II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [45]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [63] against polynomial-time adversaries. Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table IX).
III Preliminaries
In this section, we provide a concise introduction to the preliminaries. A list of notations used throughout the paper can be found in Table I.